# Deep Multimodal Learning with Missing Modality: A Survey

### v3

@@ 6.5. Robotic Vision
Discussion: Current MLMM research primarily focuses on the areas of sentiment analysis, medical diagnosis, remote sensing analysis, retrieval/captioning, and robotic vision. Among these, a significant portion of research is dedicated to addressing the missing modality problem in video, text, and audio-based sentiment analysis, MRI segmentation and clinical prediction in medical diagnosis, and multi-sensor semantic segmentation for autonomous vehicles. In contrast, research efforts in MLMM for streaming data and scientific fields are relatively limited. In addition, the majority of the publicly available datasets mentioned above are complete in terms of modalities, and naturally occurring datasets with missing modalities are rare. As a result, these tasks are often evaluated by considering all possible combinations of missing modalities based on the existing types of modalities in the dataset, and then performing training and testing on the performance of the missing modalities, supplemented by different missing modality rates. In the works we reviewed, 38% controlled for varying degrees of missing modality rates and tested accordingly, while the remaining 62% used random modality missing rates for training and testing. When categorizing missing modality rates into mild (¡30%), moderate (30%-70%), and severe (¿70%) levels, the proportions are 72.6%, 82.2%, and 69.9%, respectively. It is worth noting that a single study may perform validation across multiple missing modality rate levels. Additionally, we found that 36.4% of the works trained on incomplete training datasets (cannot access missing modality data during training), while 63.6% actually focused on testing with missing modalities. We can see that there is still small number of research on incomplete training datasets, so it is important to emphasise the need for more research to this field.
7. Open Issues and Future Research Directions
7.1. Accurate Missing Modality/Representation Data Generation
Many works (Ma et al., 2021b) have shown that recovering, reconstructing or generating missing modalities or their representations can help improve the performance of the model when encountering missing-modality samples. However, these generated modalities or representations usually have artifacts or hallucinations, which may be caused by the generative model itself or its training datasets. Therefore, exploring how to generate accurate and unbiased missing modalities or their representations will be the important research direction in the future.
7.2. Recovery or Non-Recovery Methods ?
@@
In MLMM, “to recover or not to recover the missing modality”, that is the question. According to our observations, most current approaches to handling MLMM can be broadly categorized into two types: one focusing on recovering the missing modal information so that the multimodal model can continue functioning (Wang et al., 2023a; Ma et al., 2021b), and the other attempting to make predictions using only the data modalities available (Hu et al., 2020; Zhang et al., 2024). However, some studies (Yao et al., 2024) have shown that the recovered modality information might be ignored or dominated by the existing modality information due to the imbalance missing rates of different modalities in the training set, leading the model to still rely on the available modalities when making predictions. Moreover, Lin et al. (Lin et al., 2024) have found that in situations where crucial modality information is missing, the model might be unduly influenced by less important existing modality information, resulting in incorrect outcomes. In such cases, attempting to recover missing modality information can potentially alleviate this problem. Consequently, determining under what circumstances recovery or non-recovery methods are more beneficial for model predictions, as well as how to measure whether the recovered modality information is not dominated by other modality information or is unbiased, remains a significant challenge. Addressing these issues is essential for advancing the effectiveness and reliability of deep multimodal learning with missing modality methods.
7.3. Benchmarking and Evaluations for Missing Modality Problems
@@
Benchmarking and fair evaluation play crucial roles in guiding the field. Many works are trained/tested under different missing modality settings, which makes it difficult for researchers to find works with the same settings for comparison. Also, the recent rise of large pre-trained models, such as GPT-4, has given researchers hope for achieving Artificial General Intelligence. Consequently, more large models (Liu et al., 2024a) are integrating visual, auditory, and other modal information to realize the vision of Multimodal Large Language Models. Although some research has been done based on those models, detailed analysis and benchmarks are urgently needed to evaluate the performance of MLLMs on missing modality problems. We call on the community to build a work similar to MultiBench (Liang et al., 2021), covering common datasets mentioned in Section 6 and settings of missing modality problems, to help researchers conduct more systematic research.
7.4. Method Efficiency
@@
Current MLMM methods often overlook the need for more lightweight and efficient methods. For instance, some model combinations methods (Hu et al., 2020; Xue and Marculescu, 2023) require training an independent model for each modality combination. Similarly, methods (Chen et al., 2024b; Zhang et al., 2023b) that aim to recover missing modality information typically involve using a distinct model for each kind of missing information or a large unified model to generate all modalities. Although these approaches generally perform well, they are often too heavy. In the real world, many multimodal models need to be deployed on resource-constrained devices, such as space or disaster-response robots. These devices cannot accommodate high-performance GPUs and are prone to sensor damage, which can be difficult/costly to repair. Therefore, there is an urgent need to explore efficient and lightweight solutions for MLMM that can operate effectively under these constraints.
7.5. Multimodal Streaming Data with Missing Modality
@@
Currently, only a small part of research focuses on handling missing modalities in multimodal streaming data. An example is sentiment analysis (Lin and Hu, 2023). Long sequences of multimodal data, such as RGB+X video and multimodal time series, are common in real world and essential for tasks like anomaly detection and video tracking, especially on robots. Therefore, MLMM needs to address missing modality problems in streaming data to advance this field further.
7.6. Multimodal Reinforcement Learning with Missing Modality
@@
7.7. Multimodal AI with Missing Modality for Natural Science
@@
8. Conclusion
@@
In this survey, we present the first comprehensive survey of Deep Multimodal Learning with Missing Modality. We begin with a brief introduction to the motivation of the missing modality problem and the real-world reasons that underscore its significance. Following this, we summarize the current advances based on our fine-grained taxonomy and review the applications and relevant datasets. Finally, we discuss the existing challenges and potential future directions in this field. Although more and more researchers are involved in studying the problem of missing modality, we are also concerned about some urgent issues that need to be addressed, such as a unified benchmark for testing (e.g., multimodal large language models) and the need for a wider range of applications (e.g., natural science). With our comprehensive and detailed approach, we hope that this survey will inspire more researchers to explore missing modality problems.
@@ Appendix A Paper Collection
The following shows the full names of the mentioned conference and journal abbreviations in the “Paper Collection” paragraph of Section Introduction. The collected papers come from, but are not limited to, the following conferences (such as the Association for the Advancement of Artificial Intelligence (AAAI), International Joint Conference on Artificial Intelligence (IJCAI), Conference on Neural Information Processing Systems (NeurIPS), International Conference on Learning Representations (ICLR), International Conference on Machine Learning (ICML), IEEE Conference on Computer Vision and Pattern Recognition (CVPR), International Conference on Computer Vision (ICCV), European Conference on Computer Vision (ECCV), Annual Meeting of the Association for Computational Linguistics (ACL), Conference on Empirical Methods in Natural Language Processing (EMNLP), ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD), ACM International Conference on Multimedia (ACM MM), International Conference on Medical Image Computing and Computer-Assisted Intervention (MICCAI), etc.) and journals (such as IEEE Transactions on Pattern Analysis and Machine Intelligence (TPAMI), IEEE Transactions on Image Processing (TIP), IEEE Transactions on Medical Imaging (TMI), IEEE Transactions on Multimedia (TMM), Journal of Machine Learning Research (JMLR), etc.).

### v4

@@ 6.6 Robotic Vision and Sensing
Discussion: Current MLMM research primarily focuses on the areas of affective computing, medical applications, remote sensing, information retrieval, pervasive computing and robotic vision and sensing. Among these, a significant portion of research is dedicated to addressing the missing modality problem in RGB, text, and audio-based sentiment analysis, MRI segmentation and clinical prediction, and multi-sensor segmentation for autonomous vehicles. In contrast, research efforts in MLMM for streaming data and scientific fields are relatively limited. In addition, the majority of the publicly available datasets mentioned above are complete in terms of modalities, and naturally occurring datasets with missing modalities are rare. As a result, these tasks are often evaluated by considering all possible combinations of missing modalities based on the existing types of modalities in the dataset, and then performing training and testing on the performance of the missing modalities, supplemented by different missing modality rates. Perhaps more future work could try to emulate Pervasive Computing by collecting multimodal datasets with real-world modal data from different models, including those with real sensor damage or different sampling rates of time-series information, for model validation. In the works we reviewed, 38% controlled for varying degrees of missing modality rates and tested accordingly, while the remaining 62% used random modality missing rates for training and testing. When categorizing missing modality rates into mild (<30%), moderate (30%-70%), and severe (>70%) levels, the proportions are 72.6%, 82.2%, and 69.9%, respectively. It is worth noting that a single study may perform validation across multiple missing modality rate levels. Additionally, we found that 36.4% of the works trained on incomplete training datasets (cannot access missing modality data during training), while 63.6% actually focused on testing with missing modalities. We can see that there is still small number of research on incomplete training datasets, so it is important to emphasise the need for more research to this field.
Beyond the above observations, we also note several underexplored yet crucial research gaps. First, current benchmarks rarely model realistic missingness patterns such as systematic or asynchronous sensor failures, despite their prevalence in pervasive and wearable computing. For example, in wearable activity recognition (Nweke et al., 2018), IMUs and physiological sensors frequently sample asynchronously or drop entire segments, producing structured missingness that cannot be captured by random masking. Second, the semantic importance of different modalities is typically assumed to be uniform, and there is no principled framework for measuring modality contribution or uncertainty under missing conditions. Third, evaluation protocols remain highly inconsistent across studies, with varying definitions of missing rates, train–test configurations, and dataset preprocessing. For example, SMIL (Ma et al., 2021c) considers 90% of the multimodal samples in the training set with missing modalities, but DrFuse (Yao et al., 2024) only considers the case of random missing modalities and does not consider the performance at different missing modality rates. Fourth, most existing models struggle to generalize across domains, devices, or spatiotemporal distributions, highlighting the need for scalable and transferable foundations for MLMM. Finally, there is a lack of rigorous real-world case-study evaluations and systematic testing on out-of-distribution multimodal datasets, both of which are essential for assessing robustness under practical deployment conditions. Addressing these gaps will be critical for developing MLMM systems that can reliably operate in real-world multimodal environments.
7 Open Issues and Future Research Directions
7.1 Accurate Missing Modality/Representation Data Generation
Some researchers (Ma et al., 2021c) have shown that recovering, reconstructing, or generating missing modalities or their intermediate representations can improve model robustness when encountering incomplete multimodal inputs. However, the generated content often contains artifacts or hallucinations, stemming either from the generative models themselves or from imperfections and biases in their training data. Beyond improving generative fidelity, a key future direction is to better understand informational complementarity and redundancy across modalities: not all modalities contribute equally to downstream tasks, and some provide overlapping signals that can be safely inferred, while others contain unique information that is difficult to reconstruct without introducing bias. Therefore, exploring principled ways to generate accurate, unbiased, and task-relevant missing modalities—while leveraging modality complementarity and avoiding over-reliance on redundant cues—will be an important avenue for future research.
7.2 Recovery or Non-Recovery Methods ?
@@
In MLMM, “to recover or not to recover the missing modality”, that is the question. According to our observations, most current approaches to handling MLMM can be broadly categorized into two types: one focusing on recovering the missing modal information so that the multimodal model can continue functioning (Wang et al., 2023a; Ma et al., 2021c), and the other attempting to make predictions using only the data modalities available (Hu et al., 2020; Zhang et al., 2024). However, some studies (Yao et al., 2024) have shown that the recovered modality information might be ignored or dominated by the existing modality information due to the imbalance missing rates of different modalities in the training set, leading the model to still rely on the available modalities when making predictions. Moreover, Lin et al. (2024) have found that in situations where crucial modality information is missing, the model might be unduly influenced by less important existing modality information, resulting in incorrect outcomes. In such cases, attempting to recover missing modality information can potentially alleviate this problem. Consequently, determining under what circumstances recovery or non-recovery methods are more beneficial for model predictions, as well as how to measure whether the recovered modality information is not dominated by other modality information or is unbiased, remains a significant challenge. Addressing these issues is essential for advancing the effectiveness and reliability of deep multimodal learning with missing modality methods.
7.3 Benchmarking and Evaluations for Missing Modality Problems
@@
Benchmarking and fair evaluation play crucial roles in guiding the field. Many works are trained/tested under different missing modality settings, which makes it difficult for researchers to find works with the same settings for comparison. Also, the recent rise of large pre-trained models, such as GPT-4, has given researchers hope for achieving Artificial General Intelligence. Consequently, more large models (Liu et al., 2024a) are integrating visual, auditory, and other modal information to realize the vision of Multimodal Large Language Models. It is important to note that recent multimodal models (especially transformer-based models) are typically trained on significantly larger and more diverse datasets than earlier approaches. Early studies on missing modality learning often relied on constrained, domain-specific datasets (e.g., medical imaging dataset) to validate model architectures and theoretical assumptions. In contrast, recent works leverage large-scale multimodal pretraining, sometimes involving billions of image-text or audio-visual pairs collected from the web. This expansion in data scale not only improves model generalization but also changes the nature of the learning problem, thus some recent models can implicitly infer missing information from learned priors rather than explicit modality alignment. Therefore, when comparing results across methods, it is crucial to account for these disparities in dataset scale, diversity, and accessibility, as they directly affect both performance and reproducibility. Although some research has been done based on those models, detailed analysis and benchmarks are urgently needed to evaluate the performance of MLLMs on missing modality problems. We call on the community to build a work similar to MultiBench (Liang et al., 2021), covering common datasets mentioned in section 6 and settings of missing modality problems, to help researchers conduct more systematic research.
7.4 Method Efficiency
@@
Current MLMM methods often overlook the need for more lightweight and efficient methods. For instance, some model combinations methods (Hu et al., 2020; Xue and Marculescu, 2023) require training an independent model for each modality combination. Similarly, methods (Chen et al., 2024a; Zhang et al., 2023c) that aim to recover missing modality information typically involve using a distinct model for each kind of missing information or a large unified model to generate all modalities. Although these approaches generally perform well, they are often too heavy. In the real world, many multimodal models need to be deployed on resource-constrained devices, such as space or disaster-response robots. These devices cannot accommodate high-performance GPUs and are prone to sensor damage, which can be difficult/costly to repair. Therefore, there is an urgent need to explore efficient and lightweight solutions for MLMM that can operate effectively under these constraints.
7.5 Multimodal Streaming/Temporal Data with Missing Modality
@@
Currently, only a small portion of research focuses on handling missing modalities in multimodal streaming or temporal data. An example is sentiment analysis (Lin and Hu, 2023). Long sequences of multimodal data—such as RGB+X video streams and multimodal time-series signals—are common in real-world settings and are essential for tasks like anomaly detection and video tracking, especially in robotic systems and pervasive computing. In such scenarios, missing modalities often occur asynchronously, creating not only spatial information gaps but also temporal misalignment between surviving modalities. This misalignment makes recovery, fusion, and prediction more challenging, as models must infer missing information while maintaining consistent temporal dependencies. Therefore, MLMM needs to address both missing-modality problems and temporal alignment issues in streaming data to advance this field further.
7.6 Multimodal Reinforcement Learning with Missing Modality
@@
7.7 Multimodal AI with Missing Modality for Natural Science
@@
7.8 Handling Practical Missing Modality Problems in Real-World
A promising yet still insufficiently explored direction for MLMM lies in real world scenarios—including IoT infrastructures, wearables, smart homes, mobile sensing, and distributed environmental sensor networks. Although several studies in pervasive computing have begun to address missing-modality challenges in these settings, current efforts remain fragmented and lack holistic evaluation. These environments constitute one of the most consequential real-world deployment contexts for multimodal models, where missingness is not occasional or synthetic but structural and persistent due to heterogeneous devices operating under constraints such as limited battery life, bandwidth restrictions, hardware degradation, asynchronous sampling, and intermittent connectivity. As a result, multimodal streams often exhibit long gaps, irregular intervals, and complete sensor dropout—conditions rarely captured by existing benchmarks. Empirical findings from mobile and wearable sensing further highlight the gap: human activity recognition systems frequently encounter inconsistent sampling, device failures, and user-dependent variability (Nweke et al., 2018); wrist-worn sensing studies report prolonged non-wear periods and dropout of inertial or physiological channels (Van Der Donckt et al., 2024); and systems integrating IMUs, magnetometers, and physiological sensors exhibit structural missingness induced by asynchronous fusion (Chen et al., 2023a). While these studies provide valuable insights, MLMM still lacks general frameworks capable of handling long-term partial observability, heterogeneous devices, asynchronous and irregular multimodal streams, and multi-agent inconsistencies in a unified manner. More systematic evaluations and standardized benchmarks are urgently needed. A complementary future direction involves developing resource-aware MLMM robustness strategies. Although prior work has explored lightweight methods such as signal augmentation and cross-modality distillation (Jeon et al., 2021), and energy-efficient imputation pipelines tailored for mobile health applications (Hussein et al., 2024b), these techniques remain application-specific and do not yet form a generalizable toolkit. Likewise, generative approaches such as SensorGAN (Hussein and Bhat, 2024) show promise for reconstructing missing sensor channels, but broader assessments of their scalability, latency, and reliability in real-world environments are still lacking. Overall, handling practical missing modality problems in real world presents a critical frontier for MLMM research. Despite early progress, the field still needs more comprehensive evaluations, standardized benchmarks, and general-purpose robustness principles. Ultimately, enabling multimodal AI systems that remain reliable under persistent sensor failures and dynamically evolving, heterogeneous environments will require unified methodologies and architectures explicitly designed for real-world sensing constraints.
8 Conclusion
@@
In this survey, we present the first comprehensive survey of Deep Multimodal Learning with Missing Modality. We begin with a brief introduction to the motivation of the missing modality problem and the real-world reasons that underscore its significance. Following this, we summarize the current advances based on our fine-grained taxonomy and review the applications and relevant datasets. Finally, we discuss the existing challenges and potential future directions in this field. Although more and more researchers are involved in studying the problem of missing modality, we are also concerned about some urgent issues that need to be addressed, such as a unified benchmark for testing and the need for a wider range of applications With our comprehensive and detailed approach, we hope that this survey will inspire more researchers to explore missing modality problems.
Limitations: Because our survey aims to cover deep multimodal learning with missing modalities across as many domains as possible from 2012 to 2025, our taxonomy focuses on the practical machine learning perspective rather than analysing methods in a theoretical way. Moreover, although we separately review how large models address missing modalities, the rapid development of foundation model ecosystems suggests that this taxonomy may require further refinement as new architectures, training paradigms, and modality agnostic interfaces continue to emerge.
In addition, method efficiency remains an underexplored yet increasingly critical challenge. Currently, most existing MLMM approaches rely on computationally heavy architectures, expensive cross modal alignment modules, or complex imputation pipelines. Their scalability to edge devices, resource constrained deployments, and long horizon streaming scenarios is still limited. Future research should systematically examine efficiency trade offs, including latency, memory footprint, modality specific bottlenecks, and asymmetry between training and inference, to enable practical and deployable MLMM systems.
Additionally, current benchmarks rarely capture realistic missingness patterns such as systematic, asynchronous, intermittent, or catastrophic sensor failures. Existing studies also tend to assume uniform modality importance and lack principled frameworks for quantifying modality contribution, uncertainty, or redundancy under missing conditions. Evaluation protocols remain inconsistent across works, hindering fair comparison. Beyond this, most methods still struggle to generalize across domains, devices, and spatiotemporal distributions.
Finally, while MLMM research is most active in robotic vision and sensing, medical applications, affective computing, remote sensing, and pervasive computing, its application to real-world sensing systems and natural science domains such as IoT infrastructures, wearables, smart homes, mobile sensing, distributed environmental sensor network, climate modeling, earth observation, molecular biology, and materials discovery remains sparse. These fields feature rich, heterogeneous, and often severely incomplete modalities, for example missing experimental measurements or partially observed molecular conformations. Extending MLMM research to these areas will require domain aware missingness models, physics and biology informed constraints, and better theoretical understanding.
Addressing these limitations, along with developing stronger theoretical frameworks, more realistic benchmarks, efficiency aware architectures, and broader cross domain applicability, will help guide the design of more robust, transferable, and scientifically grounded future MLMM research.
Acknowledgments
We appreciate all the valuable feedback from the anonymous reviewers and the TMLR editors. In addition, Renjie would like to thank all co-authors for their continued support.
@@ Appendix A Paper Collection
The following shows the full names of the mentioned conference and journal abbreviations in the “Paper Collection” paragraph of Section Introduction. The collected papers come from, but are not limited to, the following conferences (such as the Association for the Advancement of Artificial Intelligence (AAAI), International Joint Conference on Artificial Intelligence (IJCAI), Conference on Neural Information Processing Systems (NeurIPS), International Conference on Learning Representations (ICLR), International Conference on Machine Learning (ICML), IEEE Conference on Computer Vision and Pattern Recognition (CVPR), International Conference on Computer Vision (ICCV), European Conference on Computer Vision (ECCV), Annual Meeting of the Association for Computational Linguistics (ACL), Conference on Empirical Methods in Natural Language Processing (EMNLP), ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD), ACM International Conference on Multimedia (ACM MM), International Conference on Medical Image Computing and Computer-Assisted Intervention (MICCAI), etc.) and journals (such as IEEE Transactions on Pattern Analysis and Machine Intelligence (TPAMI), IEEE Transactions on Image Processing (TIP), IEEE Transactions on Medical Imaging (TMI), IEEE Transactions on Multimedia (TMM), Journal of Machine Learning Research (JMLR), TMLR (Transactions on Machine Learning Research), etc.).